Of the total number of spectators at a circus show, 1/4 are men. 2/5 of the remaining number of spectators are women. there are 132 women at the circus show. How many children are at the circus show?

Answer by rfer(16322)

132/(2/5)=
132*5/2=330
330-132=198 children